OpenStudy (anonymous):

Can someone help me? the problem is 5x 4/13 so it would be 5/1 x 4/13 = 20/13 = 1 7/13 how was it change to 1 7/13? please help

OpenStudy (goformit100):

@apoorvk

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[20-13=7\] there is now an extra 1, so there is the one from the mixed number and the fraction is now 7/13; you keep the denominator and the numerator changes from 20 to 7; see the simple problem above.

OpenStudy (anonymous):

Thanks @blossom1015

OpenStudy (anonymous):

okay, where is this one coming from? if you dont mind me asking.

OpenStudy (anonymous):

but again thanks a lot for helping me! =)

OpenStudy (anonymous):

oh, the one is the 13/13 that came out of the numerator.

OpenStudy (anonymous):

I see you have trouble with some basic math, no offense.

OpenStudy (anonymous):

here is a website that can help: http://www.khanacademy.org

OpenStudy (anonymous):

no yea, I know, its been a while... thanks!

OpenStudy (anonymous):

the website has practice excercises and videos which can help

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

bye, thanks again to both of ya'll =)

ganeshie8 (ganeshie8):

\[\frac{20}{13} \\ = \frac{13+7}{13} \\ = \frac{13}{13} + \frac{7}{13} \\ = 1 + \frac{7}{13} \\ = 1\frac{7}{13}\]

OpenStudy (anonymous):

thanks! ima save that =)

ganeshie8 (ganeshie8):

welcome... :)